A box contains 10,000 widgets. Pat tells you that 2% of the widgets in a box are made of copper and the


rest are made of steel. If you randomly select 15 widgets from the box, find the probability that less than 2 of the


widgets you select are made of copper.



1
Expert's answer
2021-02-24T07:46:34-0500

This is a binomial probability with n=15, p=0.02.

"P(X<2)=P(X=0)+P(X=1)=C_{15}^00.02^00.98^{15}+C_{15}^10.02^10.98^{14}="

"=0.7386+0.2261=0.9647."